# ERASE: Benchmarking Feature Selection Methods for Deep Recommender Systems

### v2

@@ 3.2. Overall Performance (RQ1)
Overall experimental results with different backbone models. In this subsection, we compare the performance of feature selection methods with different backbone models. The complete experimental results for all four backbone models are illustrated in Appendix H. For the two-stage approaches (Lasso, GBDT, RF, XGBoost, AutoField, Permutation, SHARK, and SFS), we experiment with different values of selected features during the retraining stage and adopt the best results. In the case of the single-stage approach, if it can also be modified to a two-stage method (OptFS and LPFS), we conduct experiments on both ways and record the optimal results; if it only supports single-stage (i.e., AdaFS), we directly document its results. From Table 2, we can observe the following points:
@@
  • The validity of feature selection methods is distinct across different datasets. The shallow feature selection methods perform better in Criteo and Aliccp. The possible reason is the feature value distribution in the other two datasets is very imbalanced, making shallow models prone to overfitting. Gate-based feature selection methods are good at dealing with limited data, while sensitivity-based feature selection methods achieve the best performance among all methods in datasets with rich samples. This finding can be attributed to the fact that in the data-scare scenario, gradients are too sensitive to judge feature importance. However, with sufficient data support, the stability of gradient information will be greatly improved, allowing for more accurate feature importance rankings.
@@
3.3. Effects on Number of Selection (RQ2)
Overall experimental results with different number of selected features. In this subsection, we conduct experiments on each backbone model and use the different number of selected features $K$. Figure 3 shows the effects of the number of selected features with feature selection methods. The x-axis represents the number of selected features $K$ and the y-axis represents the evaluation metrics AUC or Logloss. We only visualize the results by taking DeepFM as the backbone model here, for we find that the trends for different backbone models remain consistent. The detailed experimental results for the other backbone models are listed in Appendix I. In addition, it is worth noting that to make OptFS applicable in this experiment, we follow the drop rates of feature values in each feature field to assign importance scores to features. From Figure 3, we can find the following information:
  • The effectiveness of features selected by different methods varies in ranking when the number of selections changes. Specifically, when the number of features selected is smaller (i.e., 5 or 10), Autofield, SHARK, and SFS perform better. They obtain information through the values or gradients of gate vectors, guiding the sorting of feature importance. They are good at selecting the most informative features from the candidate feature sets. OptFS initially performs not good, as it is inherently a single-stage method. Modifying it to a two-stage method based on different feature value drop rates may not align with the goal of enhancing effectiveness through feature selection. When the number of features approaches the full set of features, the Permutation method performs well. This is because the permutation method calculates feature importance scores based on the loss of effect caused by dropping a feature from the full set. It tends to identify redundant features with less information.
@@ 3.3. Effects on Number of Selection (RQ2)
Experimental results of AUKC. To compare the robustness and stability of feature selection methods across different numbers of selections, we evaluate the selection results with the AUKC metric. Table 3 shows the experimental results of AUKC on Avazu and Criteo. We can draw the following points: 1) Overall, sensitivity-based feature selection methods outperform gate-based feature selection methods, which are superior to shallow feature selection methods. The possible reason is that gradient-based methods calculate feature importance based on partial batches, which helps prevent overfitting. The inferior performance of shallow feature selection methods may be due to the simplicity of the models, which cannot effectively capture the complex relationships between features to determine feature importance. 2) Regarding specific methods, AutoField, SHARK, and SFS exhibit the best performance, demonstrating leading effects across different datasets. It is noteworthy that AutoField significantly surpasses other gate-based feature selection methods. This may be attributed to AutoField being trained in a bi-level optimization manner, which reduces the risk of overfitting, making the selection results more stable and reliable. In conclusion, AutoField, SHARK, and SFS are superior to the other methods on the robustness and stability across different numbers of selections.
@@ 3.4. Efficiency Analysis (RQ3)
In this subsection, we answer the RQ3 from two aspects: experiments with performance limitations and with memory limitations.
@@ 3.4.2. Experiments with memory limitations
3.5. Offline Results on Industrial Dataset (RQ4)
To find whether our conslusions for feature selection methods align between large-scale industrial datasets and public datasets, we conduct comparison experiments on large-scale industrial datasets.
3.5.1. Industrial Dataset
The dataset is constructed with CVR records from 2023-11-29 to 2023-11-30 containing approximately seven hundred million samples. Within these samples, The positive sample ratio is 0.196107. The feature used in this offline dataset is aligned with the online service model, which contains 157 features (10 are dense, 36 are sparse, and 111 are multi-hot features). All of these features have been roughly proved useful in offline leave-one-out experiments.
3.5.2. Experiment
After training, we grab each method’s feature importance and perform top-k experiments. The experiment result is demonstrated in Table 6 and Figure 4. From the redundant feature elimination perspective, SHARK, AutoField, and Permutation perform the best. They eliminate about half of the feature sets (80 features) without decreasing model accuracy. However, from the effective feature mining perspective, AutoField performs best for small feature sets, aligning with our observation from the public datasets. We can also find that AutoField achieves the highest AUKC in Table 6, demonstrating its robustness and stability across different numbers of selections. LPFS’s performance is slightly misaligned with the one on the public dataset. We attribute this misalignment to the sensitive learnable polarization module.
3.6. Online Experiments (RQ5)
Specifically, we create an experiment group that reduces 30% features with our toolkit in a latency-sensitive online recommendation platform. Then, we deploy the experimental group for the A/B test with 5% traffic (approximately 1 million users). After one week’s observation, the inference latency reduces by approximately 20% while the number of video views (vv) and average play duration keeps the same. This experiment proves the effectiveness of our benchmark toolkits in the online environment.
@@ 4. Related Works
We review the related work from two perspectives, feature selection methods, as well as related benchmarks.
Feature Selection Methods. Traditional feature selection methods fall into three categories: filter, wrapper, and embedded methods (Guyon and Elisseeff, 2003). Filter methods utilize criteria to identify predictive feature fields, exemplified by the Chi2 score (Liu and Setiono, 1995) and mutual information (Vergara and Estévez, 2014). Wrapper methods employ black-box models to select predictive features, assessing the utility of feature subsets through, often, generic algorithms (Granitto et al., 2006; Shah and Kusiak, 2004). Embedded methods, on the other hand, integrate the feature selection process within the prediction model, thereby evaluating the effectiveness of feature subsets in conjunction. Notable embedded methods include LASSO (Tibshirani, 1996) and Gradient Boosting Machine (Friedman, 2001). The advent of deep learning has spurred novel approaches, such as gate-based methods, where learnable gates determine the significance of feature fields (Wang et al., 2022; Lin et al., 2022; Lyu et al., 2023; Lee et al., 2023), and sensitivity-based methods, leveraging gradient techniques to identify critical features by their sensitivity (Zhang et al., 2023; Wang et al., 2023; Fisher et al., 2019). Furthermore, some studies explored reinforcement learning to automate feature selection, using agents to pinpoint predictive features (Liu et al., 2019; Fan et al., 2020; Zhao et al., 2020).
@@
Benchmarks for Feature Selection. Most existing benchmarks are tailored for classical downstream models (Bolón-Canedo et al., 2014; Bommert et al., 2020) and rely on synthetic or domain-specific datasets (Bolón-Canedo et al., 2013; Forman, 2003; Darshan and Jaidhar, 2018). Due to their constrained relevance, these benchmarks frequently fall short in providing the necessary guidance to propel research forward in the domain of DRS feature selection. The most related work, DeepLasso (Cherepanova et al., 2023), confines its evaluation to shallow methods and a narrow set of backbone models, primarily targeting tabular learning tasks. Despite utilizing real-world datasets, the scale of these datasets pales in comparison to those encountered in DRS scenarios, thus limiting the provision of actionable insights for DRS feature selection strategies.
ERASE undertakes a thorough evaluation of shallow, gate-based, and sensitivity-based methods across an array of representative DRS backbones, leveraging both large-scale public datasets and private industrial datasets. Furthermore, ERASE proposes a novel metric designed to assess feature selection efficacy comprehensively. Standing as the inaugural comprehensive benchmark for DRS feature selection methodologies, ERASE aims to illuminate effective application strategies for these methods and to chart potential avenues for future research in this domain.
@@ Appendix A Datasets
Table 7. Dataset statistics
Dataset Avazu Criteo Movielens-1M AliCCP
Interactions 40,428,967 45,850,617 1,000,209 85,316,519
Users N/A N/A 6,040 238,635
Items N/A N/A 3,706 467,298
Interaction Type Click Click Rating (1-5) Click
Feature Num 23 39 9 23
@@
Appendix B Introduction of Feature Selection Methods
Specifically, our benchmark contains the following methods:
B.0.1. Shallow Feature Selection
  • GBDT (Friedman, 2001). Gradient-boosted decision tree (GBDT) is a popular algorithm in regression and classification tasks. GBDT achieves superior performance by continually adding trees to fit residuals. By aggregating the feature importance scores from each tree, it also serves as an effective method for feature selection.
  • RandomForest (Breiman, 2001) (RF). RandomForest is an efficient method for feature selection. It derives the feature importance by measuring how much each feature decreases the impurity in a tree, commonly using Gini impurity or entropy.
  • XGBoost (Chen and Guestrin, 2016). Extreme Gradient Boosting (XGBoost) is an efficient and scalable implementation of the gradient boosting algorithm. It ranks the importance of features by calculating the improvement of every feature on the final performance.
B.0.2. Gate-based Feature Selection
  • AutoField (Wang et al., 2022). AutoField is the first gate-based feature selection method in the deep recommender system. It designs a novel controller network to generate a 2-dimensional vector determining whether to choose this feature field or not. It is a two-stage method and selects features on the field level.
  • LPFS (Guo et al., 2022). LPFS argues that the conclusion that features with smaller weights are less important than those with larger weights may not necessarily be correct. It proposes a kind of smoothed-$l^{0}$ function that can effectively select informative features. It is a single-stage method and focuses on the field level.
B.0.3. Sensitivity-based Feature Selection
Appendix C Backbone Models
@@
Appendix D Usage Example
To enhance the credibility and reproducibility of our benchmark, we release a public repository for the convenience of research in related fields. Categorized by usage, current feature selection methods can be divided into two types: single-stage methods and two-stage methods. Detailed usage examples will be provided for these two types subsequently. In addition, in some feature selection methods, there are a few hyperparameters. To facilitate the selection of hyperparameters for optimal results, our repository integrates the automatic tuning framework NNI. With only minimal changes to the NNI configuration file, the best hyperparameter combinations can be conveniently selected.
D.1. Single-stage Methods Usage Example
For single-stage methods (e.g., AdaFS, LPFS), they don’t need the retraining procedure. In our framework, these methods can be easily implemented by the command python fs_run.py –dataset={ } –model={ } –fs={ } –train_or_search=True –retrain=False
D.2. Two-stage Methods Usage Example
For two-stage methods (e.g., Lasso, AutoField, and Permutation), if users are very familiar with the algorithm and dataset they are using and know how many features selected are best for the performance, they can simply complete the searching and retraining with a single command. python fs_run.py –dataset={ } –model={ } –fs={ } –train_or_search=True –retrain=True –k={SelectNum}
In addition, if it is very difficult to directly determine the number of feature selections $k$, then they can run the framework in the normal two-stage manner. First, run different feature selection algorithms to calculate the importance scores of different features. python fs_run.py –dataset={ } –model={ } –fs={ } –train_or_search=True –retrain=False. Then specify different $k$ values to see the impact of the number of features selected on the final result, and take the corresponding result of the optimal k as the final result. python fs_run.py –dataset={ } –model={ } –fs={ } –train_or_search=False –retrain=True –k={SelectNum} –rank_path={ } –read_feature_rank=True
D.3. Parameter Tuning Usage Example
Some feature selection algorithms require hyperparameter tuning, and manually tuning parameters is an extremely time-consuming and laborious process. Therefore, we integrate the AutoML tool NNI in the framework to help select hyperparameters. Users just need to set the parameter names and search range that need to be tuned in the configuration file of the corresponding feature selection method, and then run the NNI service with the following command. python nni_tune.py –dataset={ } –model={ } –fs={ } –train_or_search={ } –retrain={ } –k={SelectNum} –port={ServerPort}
Appendix E Metrics
Appendix F Formula Derivation of AUKC
(5) $AUKC$ $=\frac{\sum_{k=1}^{|K|}((AUC_{k}-0.5)+(AUC_{k+1}-0.5))/2}{|K|/2}$
(6) $=\frac{1}{|K|}\sum_{k=1}^{|K|}(AUC_{k}+AUC_{k+1}-1)$
(8) $=\frac{1}{|K|}\sum_{n=1}^{|N|}(AUC_{n,l}+AUC_{n,r}-1)\times\Delta l_{n}$
Appendix G Hyperparameters
Table 8. Hyperparameters for Feature Selection Methods
Method Hyperparameters
Lasso default parameters in scikit-learn
GBDT default parameters in scikit-learn
RF default parameters in scikit-learn
XGBoost default parameters in scikit-learn
AutoField update_frequency: 10
AdaFS
pretrain_epoch: 0
update_frequency: 4
OptFS gamma: 5000
LPFS
epsilon = 0.1
epsilon_update_frequency = 100
epsilon_decay = 0.9978
Permutation default parameters in scikit-learn
SHARK None
SFS num_batch_sampling = 100
As shown in Table 8, in this section, we detail the hyperparameters we utilized in this paper.
Appendix H Complete Overall Performance

### v3

@@ 3.2. Overall Performance (RQ1)
In this subsection, we compare the performance of feature selection methods with different backbone models. The complete experimental results for all four backbone models are illustrated in Appendix E. For the two-stage approaches (Lasso, GBDT, RF, XGBoost, AutoField, Permutation, SHARK, and SFS), we experiment with different values of selected features during the retraining stage and adopt the best results. In the case of the single-stage approach, if it can also be modified to a two-stage method (OptFS and LPFS), we conduct experiments on both ways and record the optimal results; if it only supports single-stage (i.e., AdaFS), we directly document its results. From Table 2, we can observe the following points:
@@
  • The validity of feature selection methods is distinct across different datasets. The shallow feature selection methods perform better in Criteo and Aliccp. The possible reason is the feature value distribution in the other two datasets is very imbalanced, making shallow models prone to overfitting. Gate-based feature selection methods are good at dealing with limited data, while sensitivity-based feature selection methods achieve the best performance among all methods in datasets with rich samples. This finding can be attributed to the fact that in the data-scarce scenario, gradients are too sensitive to judge feature importance. However, with sufficient data support, the stability of gradient information will be greatly improved, allowing for more accurate feature importance rankings.
@@
3.3. Stability Evaluation (RQ2)
In this section, we conduct experiments on each backbone model with different number of selected features $k$ to investigate their robustness and stability. Specifically, we iterate a specific list for $k$ (e.g., $[5,10,15,17,19,21]$ for Avazu) and record the corresponding performance of feature selection methods. Then, the stability of feature selection methods can be revealed from two perspectives, visual patterns of the performance curve (Figure 3) and AUKC values calculated as Equation (1) (Table 3).
In Figrue 3, the x-axis represents the number of selected features $k$ and the y-axis represents the evaluation metrics AUC or Logloss. We only visualize the results of DeepFM, for we find that the trends for different backbone models remain consistent. The detailed experimental results for the other backbone models are listed in our released repository. In addition, it is worth noting that to make OptFS applicable in this experiment, we assign importance scores to features based on the drop rates of feature values in each feature field. From Figure 3, we can find:
  • The effectiveness of features selected by different methods varies in ranking when the number of selections changes. Specifically, when the number of features selected is smaller (i.e., 5 or 10), Autofield, SHARK, and SFS perform better. They obtain information through the values or gradients of gate vectors, guiding the sorting of feature importance. They are good at selecting the most informative features from the candidate feature sets. OptFS doesn’t initially perform well, as it is inherently a single-stage method. Modifying it to a two-stage method based on different feature value drop rates may not align with the goal of enhancing effectiveness through feature selection. When the number of features approaches the full set of features, the Permutation method performs well. This is because the permutation method calculates feature importance scores based on the loss of effect caused by dropping a feature from the full set. It tends to identify redundant features with less information.
@@ 3.3. Stability Evaluation (RQ2)
To offer a more direct comparison of the stability, we evaluate the selection results using the AUKC metric. Table 3 shows AUKC results on Avazu and Criteo. We can conclude:
  • Overall, sensitivity-based feature selection methods outperform gate-based feature selection methods, both of which are superior to shallow feature selection methods. The possible reason is that gradient-based methods calculate feature importance based on partial batches, which helps to prevent overfitting. The inferior performance of shallow methods may be attributed to their excessive simplicity, which cannot effectively capture the complex relationships between features to determine feature importance.
  • Regarding specific methods, AutoField, SHARK, and SFS exhibit the best performance, demonstrating leading effects across different datasets. It is noteworthy that AutoField significantly surpasses other gate-based feature selection methods. This may be attributed to AutoField being trained in a bi-level optimization manner, which reduces the risk of overfitting, making the selection results more stable and reliable.
@@ 3.4. Efficiency Analysis (RQ3)
In this section, we answer the RQ3 from two aspects: experiments with performance limitations and with memory limitations.
@@ 3.4.2. Experiments with memory limitations
3.5. Offline Results on Industrial Dataset (RQ4)
To find whether our conclusions for feature selection methods align between large-scale industrial datasets and public datasets, we conduct comparison experiments on large-scale industrial datasets.
3.5.1. Industrial Dataset
The dataset is constructed with sampled CVR records from 2023-11-29 to 2023-11-30 containing approximately seven hundred million samples. The feature used in this offline dataset is aligned with the online service model, which contains over 157 features. All of these features have been roughly proved useful in offline leave-one-out experiments.
3.5.2. Experiment
After training, we grab each method’s feature importance and perform top-k experiments. The experiment result is demonstrated in Table 6 and Figure 4. From the redundant feature elimination perspective, SHARK, AutoField, and Permutation perform the best. They eliminate about half of the feature sets (80 features) without decreasing model accuracy. However, from an effective feature mining perspective, AutoField performs best for small feature sets, aligning with our observations from the public datasets. We can also find that AutoField achieves the highest AUKC in Table 6, demonstrating its robustness and stability across different numbers of selections. LPFS’s performance is slightly misaligned with the one on the public dataset. We attribute this misalignment to the sensitive learnable polarization module.
3.6. Online Experiments (RQ5)
Specifically, we create an experiment group that reduces 30% features with our toolkit in a latency-sensitive online recommendation platform. Then, we deploy the experimental group for the A/B test with 5% traffic (approximately 1 million users). After one week’s observation, the inference latency was reduced by approximately 20% while the number of video views (vv) and average play duration remained the same. This experiment proves the effectiveness of our benchmark toolkits in the online environment.
3.7. Similarity Visualization (RQ6)
Figure 5. Similarity between feature selection methods.
In this section, we visualize the similarity between feature rankings of different feature selection methods in DRS across 4 public datasets. Figure 5 shows the heatmap of pair-wise Spearman correlation similarities. In the heatmap, the x-axis and y-axis represent different feature selection methods. The color of each square indicates the level of similarity. The more the color leans towards dark green, the more similar between the feature rankings derived from the two feature selection methods. Conversely, the more the color leans towards light green, the greater the difference in the feature rankings selected by the two methods.
@@ 4. Related Works
Feature Selection Methods. Machine learning has shown superior results in data mining (Jia et al., 2024b; Zhang et al., 2024; Li et al., 2023a; Jia et al., 2024c; Karpathy and Fei-Fei, 2015; Jia et al., 2023, 2024a). The modeling performance relies heavily on input features, making feature selection a crucial part of feature engineering. Traditional feature selection methods fall into three categories: filter, wrapper, and embedded methods (Guyon and Elisseeff, 2003). Filter methods utilize criteria to identify predictive feature fields, exemplified by the Chi2 score (Liu and Setiono, 1995) and mutual information (Vergara and Estévez, 2014). Wrapper methods employ black-box models to select predictive features, assessing the utility of feature subsets through, often, generic algorithms (Granitto et al., 2006; Shah and Kusiak, 2004). Embedded methods, on the other hand, integrate the feature selection process within the prediction model, thereby evaluating the effectiveness of feature subsets in conjunction. Notable embedded methods include LASSO (Tibshirani, 1996) and Gradient Boosting Machine (Friedman, 2001). The advent of deep learning has spurred novel approaches (Liu et al., 2024; Chen et al., 2024; Zhaok et al., 2021; Zhao et al., 2021b; Liu et al., 2020; Zhao et al., 2021a; Li et al., 2023b), such as gate-based methods, where learnable gates determine the significance of feature fields (Wang et al., 2022; Lin et al., 2022; Lyu et al., 2023; Lee et al., 2023), and sensitivity-based methods, leveraging gradient techniques to identify critical features by their sensitivity (Zhang et al., 2023; Wang et al., 2023; Fisher et al., 2019). Furthermore, some studies explored reinforcement learning to automate feature selection, using agents to pinpoint predictive features (Liu et al., 2019; Fan et al., 2020; Zhao et al., 2020).
@@
Benchmarks for Feature Selection. Most existing benchmarks are tailored for classical downstream models (Bolón-Canedo et al., 2014; Bommert et al., 2020) and rely on synthetic or domain-specific datasets (Bolón-Canedo et al., 2013; Forman, 2003; Darshan and Jaidhar, 2018). Due to their constrained relevance, these benchmarks frequently fall short in providing the necessary guidance to propel research forward in the domain of DRS feature selection. The most related work, DeepLasso (Cherepanova et al., 2023), confines its evaluation to shallow methods and a narrow set of backbone models, primarily targeting tabular learning tasks. Despite utilizing real-world datasets, the scale of these datasets pales in comparison to those encountered in DRS scenarios, thus limiting the provision of actionable insights for DRS feature selection strategies. ERASE undertakes a thorough evaluation of shallow, gate-based, and sensitivity-based methods across an array of representative DRS backbones, leveraging both large-scale public datasets and private industrial datasets. Furthermore, ERASE proposes a novel metric designed to assess feature selection efficacy comprehensively.
@@ 5. Conclusion
6. Acknowledgments
This research was partially supported by Huawei (Huawei Innovation Research Program), Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No. ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), and SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046, No.7020074).
@@ Appendix A Datasets
Appendix B Backbone Models